OpenStudy (anonymous):

In a right triangle the cosine of an acute angle is 1/2 and the hypotenuse measures 7 inches. What is the length of the shortest side of the triangle?

OpenStudy (lgbasallote):

cosine is opposite over hypotenuse so opposite over hypotenuse is 1/2 however it says that hypotenuse should be 7..so you put it in terms of 7 1/2 = x/7 x = 3.5 the shortest side would be \[\sqrt{7^2 - (3.5)^2}\] got it?

OpenStudy (lgbasallote):

hmm wait...i tried it and it turns out that the shortest side will be 3.5 because \(\sqrt{7^2 - (3.5)^2} = 6.06\)

OpenStudy (quint4698):

which one right